OpenStudy (angelwings996):

what is the solution to the equation? (-9x+5)^1/2=(3+4x)^1/2

OpenStudy (anonymous):

x=2/13

OpenStudy (anonymous):

\[\sqrt{-9x+5}=\sqrt{3+4x}\]

OpenStudy (anonymous):

\[(\sqrt{-9x+5})^2=(\sqrt{3+4x})^2\]

OpenStudy (angelwings996):

Is 2/13 right ? or no

OpenStudy (anonymous):

\[-9x+5=3+4x\]

OpenStudy (anonymous):

is r8

OpenStudy (anonymous):

\[5-3=9x+4x\]

OpenStudy (anonymous):

\[2=13x\]

OpenStudy (anonymous):

divd by 13 both ll be \[2/13=x\]

OpenStudy (angelwings996):

Okay thanks :)

OpenStudy (anonymous):

ur welcome